# Hierarchical Scheduling for Multi-Vector Image Retrieval

### v1

@@
Abstract
To effectively leverage user-specific data, retrieval-augmented generation (RAG) is employed in multimodal large language model (MLLM) applications. However, conventional retrieval approaches often suffer from limited retrieval accuracy. Recent advances in multi-vector retrieval (MVR) improve accuracy by decomposing queries and matching against segmented images. They still suffer from sub-optimal accuracy and efficiency, overlooking alignment between the query and varying image objects and redundant fine-grained image segments. In this work, we present an efficient scheduling framework for image retrieval – HiMIR. First, we introduce a novel hierarchical paradigm, employing multiple intermediate granularities for varying image objects to enhance alignment. Second, we minimize redundancy in retrieval by leveraging cross-hierarchy similarity consistency and hierarchy sparsity to minimize unnecessary matching computation. Furthermore, we configure parameters for each dataset automatically for practicality across diverse scenarios. Our empirical study shows that, HiMIR not only achieves substantial accuracy improvements but also reduces computation by up to $3.5\times$ over the existing MVR system.
footnotetext: $\dagger$ Equal Contribution. * Corresponding to Xiang Chen (xiang.chen@pku.edu.cn).
I Introduction
The rapid advancement of large language models (LLMs) has enabled emerging applications such as intelligent agents and personal assistants [1]. However, LLMs are not inherently capable of effectively leveraging user-specific data. To address this limitation, retrieval-augmented generation (RAG) has been widely adopted, where external data relevant to a user query is retrieved and incorporated into the generation process to improve accuracy and relevance [2]. Meanwhile, when RAG is extended to multimodal LLM systems, specifically with image retrieval, it computes the similarity between the feature vectors of language prompts and particular image objects [3]. Most multimodal RAG systems in production, however, follow a simple one-shot paradigm. As shown in Fig. 1, MVR embeds an entire query and an entire image into a single global vector, referred to as “1 Mode” in this context. While efficient, this single-vector retrieval inevitably loses fine-grained object information, leading to unsatisfactory retrieval accuracy for complex or semantically diverse image content.
@@ I Introduction
Figure 1: Multimodal Image Retrieval
However, despite its accuracy benefits, MVR introduces several challenges in terms of efficiency and practicality. First, the number of decomposed vectors (N) is often set to be large in order to capture fine-grained objects. Yet, selecting the optimal N is highly non-trivial: too small an N fails to fully represent image granularity, while too large an N breaks the integrity of objects. Second, finer decomposition inherently amplifies computation complexity, as each additional image segment takes extra similarity calculation against multiple sub-queries. Finally, although finer-grained decomposition reveals structural properties such as redundancy across image segments and sparsity within query–image alignments, prior work has largely overlooked exploiting these opportunities [4, 5].
Figure 2: Image Retrieval with Hierarchical Decomposition
These challenges motivate the design of HiMIR, a retrieval scheduling framework that introduces hierarchical decomposition into multi-vector image retrieval:
From the algorithm perspective, unlike the conventional “1+N Mode” MVR, HiMIR extends it into a “1+M+N Mode” as shown in Fig. 2, where M represents a series of intermediate image segmentation granularities and hierarchical query matching. This hierarchy effectively adapts to images with varying object granularities, significantly improving alignment robustness and, therefore, retrieval accuracy.
From the computing perspective, although “1+M+N Mode” seems to further increase the computation complexity compared to “1+N” MVR, it actually exposes a more favorable computing optimization space. It exploits the cross-hierarchy consistency of retrieval information to reduce redundant matching, while simultaneously detecting and pruning sparsity per-hierarchy to minimize unnecessary computations.
From the design automation perspective, HiMIR further enhances practicality by supporting automatic configuration across datasets. Given the dataset variability, it conducts a lightweight profiling of dataset characteristics and derives the optimal parameterization for hierarchical decomposition. This enables joint automation of both algorithmic decomposition and computational scheduling, significantly improving adaptability across diverse deployment scenarios.
Following such a design methodology of HiMIR, this work made the following contributions:
  • We present a hierarchical decomposition framework for multi-vector image retrieval, offering a novel approach adaptive to diverse image granularities.
  • We systematically exploit sparsity in multi-vector retrieval, enabling a runtime acceleration mechanism that eliminates redundant computations.
  • We integrate these techniques into an automated framework that jointly optimizes accuracy and efficiency, enabling robust and adaptive deployment across scenarios.
@@
II Background
II-A Decomposition in Retrieval
In multimodal RAG systems, the retrieval task aims to identify the most relevant image for a given input prompt query. Both the query and images are embedded into a shared vector space using text–image embedding models [6, 7], and the top-$K$ images with the highest similarity scores to the query are retrieved. The retrieval process can be expressed as Eq. 1, where $Q$ denotes the query, $D$ the image set, and $E(\cdot)$ the embedding model. $\operatorname{SIM}$ represents a similarity operator (e.g., cosine similarity, dot product, or L1- distance).
$\underset{0<i\leq N_{D}}{\operatorname{TopK}}(\operatorname{Score}(Q,D_{i}))=\operatorname{TopK}(\operatorname{SIM}(E(Q),E(D_{i})))$ (1)
However, encoding a semantically complex query or image into a single vector inevitably incurs information loss, resulting in sub-optimal retrieval quality. To address this issue, one-shot decomposition-based methods, or MVR (e.g., ColBERT and its variants [8, 9]), have been proposed. As formulated in Eq. 2, MVR decomposes a query into multiple sub-queries $q_{i}$} and each image into multiple segments $D_{i,j}$}. The overall score is computed as the product of the scores of all sub-queries, where the score of each sub-query is defined as the maximum similarity against all image segments. In other words, each sub-query matches its most relevant image segment, ensuring that all semantic components of the query are satisfied. To further improve retrieval accuracy, recent studies [4] adopt the “1+N Mode”, which aggregates scores computed with and without decomposition.
$\begin{split}\operatorname{Score}(Q,D_{i})=&\operatorname{SIM}(E(Q),E(D_{i}))\\ +&\prod_{i=1}^{N_{q}}\max_{1\leq j\leq m}\operatorname{SIM}(E(q_{i}),E(D_{i,j})).\end{split}$ (2)
@@ II-B Decomposition Granularity
Query decomposition granularity. A query should be decomposed into semantically independent sub-queries to achieve precise alignment with image objects. Coarse-grained decomposition often results in information loss, while overly fine-grained ones like token-level [8], may break semantic integrity and cause spurious matches. Recent work [4] leverages fine-tuned LLMs [10] to adaptively decompose complex queries.
@@
Image decomposition granularity. Unlike text retrieval, where segmentation boundaries naturally exist at the sentence or paragraph level, image retrieval lacks inherent structural boundaries. Consequently, fast adaptive segmentation methods such as SLIC (Simple Linear Iterative Clustering) [11] are employed to partition images into segments, with granularity typically set empirically in the range of 4 to 64.
The major limitation of existing image decomposition approaches is that granularity is empirically pre-defined and lacks runtime adaptability.
II-C Approximate Retrieval Acceleration
@@
As the complexity and scale of retrieval grow, computational redundancy becomes non-trivial, motivating the adoption of approximate methods that trade accuracy for efficiency. Most existing approximate acceleration techniques are algorithm-agnostic: indexing-based methods, such as IVF[12], HNSW [13], and PLAID [14], reduce the number of retrieval candidates for each query via clustering embedding vectors; quantization-based approaches such as IVF-PQ [12] reduce similarity computation latency by lowering data bit-width, thereby enhancing hardware parallelism. While, our approach exploits redundancy inherent in the MVR algorithm itself, which is largely orthogonal to these techniques and can thus be seamlessly integrated into existing optimizations.
III Hierarchical Decomposition
@@
In this section, we present the analysis for the mis-alignment between query and image segments regarding retrieval granularities. And we propose a novel hierarchical decomposition algorithm framework, which both enables better alignment and exposes new scheduling opportunities.
Figure 3: Decomposition Granularity and Accuracy Impact
III-A Granularity Misalignment Analysis
@@
Though the “1+N Mode” achieves non-trivial accuracy improvements, it still suffers from misalignment between fixed decomposition granularity and the variable scales of image objects. As shown in Fig. 3(a), which visualizes an image (same as in Fig. 5) decomposed into 25 segments with ground-truth objects marked by colored rectangles, such misalignment becomes evident. For certain objects, such as the corkboard (highlighted in red), decomposition is beneficial: irrelevant regions are cropped out while the main body of the object remains intact. However, other objects are adversely affected when the granularity is misaligned: objects may be fragmented into parts (e.g., the shelves marked in green) or merged with irrelevant regions (e.g., the toy marked in yellow).
@@
To further prove this observation, we conduct a preliminary experiment by profiling accuracy across a set of granularities using the “1+N Mode” in Eq. 2. The results on two datasets are plotted with a solid line in Fig. 3(b). As the decomposition granularity becomes finer, the retrieval accuracy generally shows an upward trend and then levels off. However, accuracy decreases at both certain granularities in the middle and the finest ones. This can be attributed to the misalignment. Hence, to attain higher retrieval accuracy, we need to align the decomposition granularity “N” with each object. This could hardly be implemented with a single fixed granularity, so we turn to the aggregation of multiple granularities.
III-B Hierarchical Aggregation
@@
Inspired by the analysis above, we want to align each required object with its best-match granularity for accuracy improvement. Thus, we extend the “1+N Mode” MVR algorithm by building the hierarchy of granularities, and searching along it for the most suitable granualrity of each sub-query (corresponding to an image object). The proposed algorithm can be denoted “1+M+N Mode” as illustrated in Fig. 2. Specifically, the images are segmented hierarchically into multiple granularities, each denoting a different patch size. The similarity score between each sub-query and image segments at a specific granularity is computed in the same way as Eq. 2. The final score for each sub-query is aggregated by selecting the maximum similarity score among all granularities, which represents the best alignment. With granularity denoted $g$, segment count in granularity $g$ denoted $N_{g}$, and the total number of granularities denoted $N_{G}$, we can extend Eq. 2 into Eq. 3:
$\operatorname{Score}(Q,D_{i})=\prod_{k=1}^{N_{q}}\max_{g=1}^{N_{G}}\max_{j=1}^{N_{g}}\operatorname{SIM}(E(q_{k}),E(D_{i,j}^{g}))$ (3)
@@
Our “1+M+N Mode” is superior to the prevalent “1+N Mode” in two aspects. First, by considering all possible granularities, it adaptively leverages the most suitable segment size for each sub-query. As shown in Fig. 5, the sub-query “toy” matches the granularity of 9 while the other two objects match the granularity of 16. As shown in Fig. 3, the accuracy of our“1+M+N Mode” is always higher than “1+N Mode” and grows monotonically as the granularity becomes finer. Furthermore, our approach provides flexibility for an accuracy-performance trade-off by allowing for computing the relevance score on a selected subset of the hierarchy, which will be elaborated in the next section.
IV Computing Efficiency Exploration
@@
The proposed hierarchical framework decouples the retrieval process into structured levels, enabling systematic analysis of information redundancy and principled optimization of computational efficiency.
Figure 4: Computing Redundancy Analysis
IV-A Redundancy Analysis with Retrieval Granularity
Figure 5: Computation Optimization in HiMIR
@@ IV-A Redundancy Analysis with Retrieval Granularity
Deriving from the computation in Eq. 3, redundancy can be analyzed from two perspectives: the image dimension $D_{i}$ and the granularity dimension $D^{g}$.
@@
Image dimension. Intuitively, coarse-grained image segments still preserve information about finer-grained objects, despite embedding loss. Neglecting the reuse of such information, which is consistent across the granularity hierarchy, leads to significant redundant computation. We validate this hypothesis by evaluating the top-$k$ recall and the rank distribution of ground-truth images across granularities, as shown in Fig. 4(a). The results indicate that ground-truth images consistently appear near the top ranks at all granularities. This observation suggests that even coarse-grained representations are sufficient to separate relevant images from irrelevant ones.
Granularity dimension. Although object sizes vary across images, they are not evenly distributed across all granularities. Exhaustively scanning every granularity for each query, therefore, introduces redundancy. To quantify this, we compute the best-matching granularity for each sub-query, and the distribution is illustrated in Fig. 4(b). The results reveal that certain granularities are rarely utilized, leading to unnecessary similarity computations for all queries. Moreover, the finest-grained representation is not always required, which further contributes to wasted computation in some cases.
@@
Based on our analysis of granularity and redundancy, we identify three complementary dimensions of computational optimization within hierarchical decomposition. As illustrated in Fig. 5, these dimensions highlight distinct opportunities for reducing redundant computations and improving efficiency, which we elaborate on in the following subsections.
@@
Image granularity analysis shows that the similarity score of ground-truth image consistently ranks near the top across all granularities. As more granularities are considered, the rank of the ground-truth image improves, whereas images ranked lower rarely benefit, even with finest-grained matching. This cross-hierarchy consistency suggests that performing fine-grained retrieval on low-ranking images is largely redundant. Based on this observation, we design an online optimization algorithm that tracks the score ranking of images in each granularity hierarchy. By eliminating costly fine-grained computations on low-confidence candidate images, the algorithm reduces runtime overhead without compromising accuracy.

### v2

# MIRAGE: Runtime Scheduling for Multi-Vector Image Retrieval with Hierarchical Decomposition

@@
Abstract.
To effectively leverage user-specific data, retrieval-augmented generation (RAG) is employed in multimodal large language model (MLLM) applications. However, conventional retrieval approaches often suffer from limited retrieval accuracy. Recent advances in multi-vector retrieval (MVR) improve accuracy by decomposing queries and matching against segmented images. They still suffer from sub-optimal accuracy and efficiency, overlooking alignment between the query and varying image objects and redundant fine-grained image segments. In this work, we present an efficient scheduling framework for image retrieval – MIRAGE. First, we introduce a novel hierarchical paradigm, employing multiple intermediate granularities for varying image objects to enhance alignment. Second, we minimize the overhead of retrieval by leveraging cross-hierarchy ranking consistency and hierarchy redundancy to eliminate unnecessary matching computation. Furthermore, we configure parameters for each dataset automatically for practicality across diverse scenarios. Our empirical study shows that, MIRAGE not only achieves substantial accuracy improvements but also reduces computation by up to $3.5\times$ over the existing MVR system.
1. Introduction
The rapid advancement of large language models (LLMs) has enabled emerging applications such as intelligent agents and personal assistants (li2024personal). However, LLMs are not inherently capable of effectively leveraging user-specific data. To address this limitation, retrieval-augmented generation (RAG) has been widely adopted, where external data relevant to a user query is retrieved and incorporated into the generation process to improve accuracy and relevance (zhao2024ragsurvey). Meanwhile, when RAG is extended to multimodal LLM systems, specifically with image retrieval, it computes the similarity between the feature vectors of language prompts and particular image objects (abootorabi2025askmodality). Most multimodal RAG systems in production, however, follow a simple one-shot paradigm. As shown in Fig. 1, MVR embeds an entire query and an entire image into a single global vector, referred to as “1 Mode” in this context. While efficient, this single-vector retrieval inevitably loses fine-grained object information, leading to unsatisfactory retrieval accuracy for complex or semantically diverse image content.
@@ 1. Introduction
Figure 1. Multimodal Image Retrieval
A AA
However, despite its accuracy benefits, MVR introduces several challenges in terms of efficiency and practicality. First, the number of decomposed vectors (N) is often set to be large in order to capture fine-grained objects. Yet, selecting the optimal N is highly non-trivial: small N fails to fully represent image granularity, while large N breaks the integrity of objects. Second, finer decomposition inherently amplifies computation complexity, as each additional image segment takes extra similarity calculation against multiple sub-queries. Finally, although finer-grained decomposition reveals structural properties such as information consistency across image segments and redundancy within query–image alignments, prior work has largely overlooked exploiting these opportunities (liu2025poqd; aiger2025globaltolocal).
Figure 2. Image Retrieval with Hierarchical Decomposition
A AA
These challenges motivate the design of MIRAGE, a retrieval scheduling framework that introduces hierarchical decomposition into multi-vector image retrieval:
From the algorithm perspective, unlike the conventional “1+N Mode” MVR, MIRAGE extends it into a “1+M+N Mode” as shown in Fig. 2, where M represents a series of intermediate image segmentation granularities and hierarchical query matching. This hierarchy effectively adapts to images with varying object scales, significantly improving alignment robustness and, therefore, retrieval accuracy.
From the computing perspective, although “1+M+N Mode” seems to further increase the computation complexity compared to “1+N” MVR, it actually exposes a more favorable computing optimization space. It exploits the cross-hierarchy consistency of retrieval information to reduce redundant matching, while detecting and pruning unnecessary hierarchies to minimize computational overhead.
From the design automation perspective, MIRAGE further enhances practicality by supporting automated configuration across datasets. Given the dataset variation, it conducts a lightweight profiling of the dataset and derives the optimal parameterization for hierarchical decomposition. This enables joint automation of algorithmic decomposition and computational scheduling, significantly improving adaptability across diverse deployment scenarios.
Following such a design methodology of MIRAGE, this work made the following contributions: (1) We present a hierarchical decomposition framework for multi-vector image retrieval, offering a novel approach adaptive to diverse image granularities. (2) We systematically exploit redundancy in multi-vector retrieval, enabling runtime acceleration mechanisms that eliminate unnecessary computations. (3) We integrate these techniques into an automated framework that jointly optimizes accuracy and efficiency, enabling robust and adaptive deployment across scenarios.
@@
2. Background
2.1. Decomposition in Retrieval
In multimodal RAG systems, the retrieval task aims to identify the most relevant image for a given input prompt query. Both the query and images are embedded into a shared vector space using text–image embedding models (radford2021learning; li2023bge1), and the top-$K$ images with the highest similarity scores to the query are retrieved. The retrieval process can be expressed as Eq. 1, where $Q$ denotes the query, $D$ the image set, and $E(\cdot)$ the embedding model. $\operatorname{SIM}$ represents a similarity operator (e.g., cosine similarity or dot product).
(1) $\underset{0<i\leq N_{D}}{\operatorname{TopK}}(\operatorname{Score}(Q,D_{i}))=\operatorname{TopK}(\operatorname{SIM}(E(Q),E(D_{i})))$
However, encoding a semantically complex query or image into a single vector inevitably incurs information loss, resulting in sub-optimal accuracy. To address this issue, one-shot decomposition-based methods, or MVR (e.g., ColBERT and its variants (khattab2020colbert; santhanam2022colbertv2)), have been proposed. As formulated in Eq. 2, MVR decomposes a query into multiple sub-queries $q_{i}$} and each image into multiple segments $D_{i,j}$}. The overall score is computed as the product of the scores of all sub-queries, where the score of each sub-query is defined as the maximum similarity against all image segments. In other words, each sub-query matches its most relevant image segment, ensuring that all semantic components of the query are satisfied. To further improve retrieval accuracy, recent studies (liu2025poqd) adopt the “1+N Mode”, which aggregates scores computed with and without decomposition.
(2) $\begin{split}\operatorname{Score}(Q,D_{i})&=\operatorname{SIM}(E(Q),E(D_{i}))\\ &+\prod_{k=1}^{N_{q}}\max\limits_{j=1}^{N_{g}}\operatorname{SIM}(E(q_{k}),E(D_{i,j})).\end{split}$
@@ 2.2. Decomposition Granularity
Query decomposition granularity. A query should be decomposed into semantically independent sub-queries to achieve precise alignment with image objects. Coarse-grained decomposition often results in information loss, while overly fine-grained ones, like token-level (khattab2020colbert), may break semantic integrity and cause spurious matches. Recent work (liu2025poqd) leverages fine-tuned LLMs (grattafiori2024llama3herdmodels) to adaptively decompose complex queries.
@@
Image decomposition granularity. Unlike text retrieval, where segmentation boundaries naturally exist at the sentence or paragraph level, image retrieval lacks inherent structural boundaries. Consequently, fast adaptive segmentation methods such as SLIC (Simple Linear Iterative Clustering) (achanta2012slic) are employed to partition images into segments. The major limitation of existing image decomposition approaches is that granularity is pre-defined and lacks runtime adaptability.
2.3. Approximate Retrieval Acceleration
@@
As the complexity and scale of retrieval grow, computational redundancy becomes non-trivial, motivating the adoption of approximate methods that trade accuracy for efficiency. Most existing approximate acceleration techniques are algorithm-agnostic: Indexing-based methods, such as IVF(douze2024faiss), HNSW (malkov2018approximate), and PLAID (santhanam2022plaid), reduce the number of retrieval candidates for each query via clustering embedding vectors; Quantization-based approaches such as IVF-PQ (douze2024faiss) reduce similarity computation latency by lowering data bit-width, thereby enhancing hardware parallelism. While, our approach exploits redundancy inherent in the MVR algorithm itself, which is largely orthogonal to these techniques and can thus be seamlessly integrated into existing optimizations.
3. Hierarchical Decomposition
@@
In this section, we analyze the misalignment between queries and image segments regarding retrieval granularities. We then propose a hierarchical decomposition framework, which enables better alignment and exposes new scheduling opportunities.
3.1. Granularity Misalignment Analysis
@@
Though the “1+N Mode” achieves non-trivial accuracy improvements, it still suffers from misalignment between fixed decomposition granularity and the variable scales of image objects. As shown in Fig. 3(a), where the image is decomposed into 25 segments with ground-truth objects highlighted by colored rectangles, the misalignment becomes evident. For certain objects, such as the corkboard (highlighted in red), decomposition is beneficial: irrelevant regions are cropped out while the main body of the object remains intact. However, other objects are adversely affected when the granularity is misaligned: objects may be fragmented into parts (e.g., the shelves marked in green) or merged with irrelevant regions (e.g., the toy marked in yellow).
@@
To further prove this observation, we conduct a preliminary experiment by profiling accuracy across a set of granularities using the “1+N Mode” in Eq. 2. The results on two datasets are plotted with solid lines in Fig. 3(b). As the decomposition granularity becomes finer, the retrieval accuracy generally shows an upward trend and then levels off. However, accuracy fluctuates at both certain granularities in the middle and the finest ones. This can be attributed to the misalignment. Hence, to achieve higher retrieval accuracy, the decomposition granularity “N” should ideally align with the scale of each object, which is infeasible with a single fixed granularity. We therefore turn to the aggregation of multiple granularities.
Figure 3. Decomposition Hierarchy and Accuracy Impact
A AA
3.2. Hierarchical Aggregation
@@
Inspired by the analysis above, we want to align each object with its best-matching granularity to improve retrieval accuracy. Thus, we extend the “1+N Mode” by constructing the hierarchy of granularities and searching along it for the most suitable granularity of each sub-query. The proposed algorithm can be denoted as “1+M+N Mode” as illustrated in Fig. 2. Specifically, each image is hierarchically segmented into multiple granularities, each corresponding to a different patch size. We then iterate from coarse to fine granularities, computing the similarity score between each sub-query and image segments at every granularity as in Eq. 2. After each iteration, the score of each sub-query is updated by selecting the maximum similarity score among all examined granularities, representing its best alignment. With granularity denoted $g$, segment count in granularity $g$ denoted $N_{g}$, and the total number of granularity hierarchies denoted $N_{G}$, we can extend Eq. 2 into Eq. 3:
(3) $\begin{split}\operatorname{Score}(Q,D_{i})=&\operatorname{SIM}(E(Q),E(D_{i}))\\ +&\prod_{k=1}^{N_{q}}\max_{g=1}^{N_{G}}\max_{j=1}^{N_{g}}\operatorname{SIM}(E(q_{k}),E(D_{i,j}^{g}))\end{split}$
@@
Our “1+M+N Mode” is superior to the prevalent “1+N Mode” in two aspects. First, by considering all possible granularities, it adaptively leverages the most suitable segment size for each sub-query. As shown in Fig. 2, the sub-query “chairs” matches the granularity of 4 while the sub-queries “computer” and “keyboard” match the granularity of 16. As shown in Fig. 3, the accuracy of our“1+M+N Mode”—which aggregates the scores of each sub-query across all prefix granularities—is always higher than “1+N Mode” and grows monotonically as the granularity becomes finer. Furthermore, our approach provides flexibility for an accuracy-performance trade-off by allowing for computing the relevance score on a selected subset of the hierarchy, which will be elaborated in the next section.
4. Computational Efficiency Exploration
@@
The proposed hierarchical framework decouples the retrieval process into structured levels, enabling systematic analysis of information redundancy and optimization of computational efficiency.
Figure 4. Redundancy Analysis: (a) Similarity of ground-truth image consistently ranks near the top across hierarchies. (b) The deepest hierarchies are not always needed. (c) Removing hollow hierarchies causes trivial accuracy loss.
A AA
4.1. Computational Redundancy Analysis
Figure 5. Computational Optimization in MIRAGE
A AA
@@ 4.1. Computational Redundancy Analysis
Deriving from Eq. 3, computation redundancy can be attributed to two dimensions: image $D_{i}$ and granularity $D^{g}$.
@@
Low-Similarity Image Long-Tail. Intuitively, coarse-grained image segments still preserve information about finer-grained objects, despite embedding loss. We argue that the coarse-grained similarity scores obtained in the early iterations are already sufficient to distinguish relevant images from irrelevant ones. We validate this hypothesis by evaluating the top-$k$ recall rate of ground-truth at different hierarchy depths. Fig. 4(a) shows that the ground-truth consistently ranks near the top. This suggests that further computing costly fine-grained similarity is redundant for a large portion of low-similarity images, which forms a long tail as shown in Fig. 5(a).
Unnecessary Deep Traversal. Although deeper hierarchies capture finer-grained objects, we observe that not all queries align with the finest-grained representations. According to Eq. 3, the similarity score between a query and an image stabilizes once all the sub-queries have found their highest-similarity patch. Therefore, exhaustively traversing all hierarchies for each query introduces substantial redundancy. To quantify this effect, we measure the distribution of (i) the granularity of the best-matching patch for each sub-query, and (ii) the deepest hierarchy at which the similarity score between the full query and an image converges, as shown in Fig. 4(b). The results indicate that deep hierarchies are not always necessary for every query, offering optimization opportunities.
Hollow Hierarchies. When adjacent hierarchies differ only slightly, an object captured by an intermediate hierarchy is also likely to be captured by neighboring hierarchies. This results in redundant intermediate hierarchies, which are referred to as “hollows”. Fig. 5(c) gives an example where both people and the sign are captured by granularity 9, yet higher-similarity matches exist at adjacent hierarchies. To evaluate this effect, we start from the granularity set with a stride of 8 and iteratively remove intermediate granularities until the final set has a stride of 16. The results in Fig. 4(c) show that removing these redundancy hierarchies has minimal impact on accuracy.
@@
Based on our analysis of computational redundancy, we identify three complementary dimensions of computational optimization within hierarchical decomposition. As illustrated in Fig. 5, these dimensions highlight distinct opportunities for efficiency improvement, which we elaborate on in the following subsections.
Algorithm 1 Online Redundancy Reduction
1:procedure ProcessQuery($Q$, $\{q_{k}\}$)
2:   Let $\operatorname{Score}[N_{D}]$, $\operatorname{SubScore}[N_{D},N_{q}]$, $\operatorname{TopK}[N_{G}]$
3:   for $g:=1$ to $N_{G}$ do
4:   for $i:=1$ to $N_{D}\cdot t_{g-1}$ do
5:    for $k:=1$ to $N_{q}$, $j:=1$ to $N_{g}$ do
6:     $\operatorname{SubScore}_{i,k}:=\max(\operatorname{SIM}(E(q_{k}),E(D_{i,j}^{g})),\operatorname{SubScore}_{i,k})$     $\operatorname{Score}_{i}:=\operatorname{SIM}(E(Q),E(D_{i}))+\prod_{k=1}^{N_{q}}\operatorname{SubScore}_{i,k}$
@@
8:   $\operatorname{TopK}[g]:=D[1:N_{K}]$
9:   if $\operatorname{K\tau}(\text{TopK}[g],\text{TopK}[g-1])\geq\tau$ then
10:    break
11:   return $\operatorname{TopK}[g]$
4.2. Low-similarity Tail Pruning
Fig. 4(a) shows that the scores of ground-truth images consistently appear among the top-ranked images even with coarse granularities. As finer granularities are introduced in deeper hierarchy levels, the top-ranked results become more precise; however, the cost of similarity computation increases substantially due to the expanded number of image segments. This motivates pruning low-ranked images from subsequent ranking updates to avoid redundant computation. We propose an online optimization algorithm that tracks image rankings across iterations. By pruning expensive fine-grained evaluations for low-similarity candidates, the algorithm effectively reduces runtime overhead without sacrificing accuracy.